A number when divided by a divisor left remainder 23.When twice of the number was divided by the same divisor , remainder was 11.Find the divisor?
Answer by ikleyn(52786)

.
A number when divided by a divisor left remainder 23.When twice of the number was divided by the same divisor , remainder was 11.Find the divisor?
~~~~~~~~~~~~~~~~~~~~~~~~~~

2*23 - 11 = 46 - 11 = 35.